OpenStudy (anonymous):

Find the x-intercept of the tangent line to the graph of f(x) = 2x − 3 cos(x) at the point (0, f(0)).

OpenStudy (anonymous):

number four

OpenStudy (anonymous):

\[f(x)=y=2x-3\cos(x)\] \[y^{'}=2+3\sin(x)\]Thus (0,f(0))=(0,-3) and\[f^{'}(0)=2\] Then the equation of the tangent line is \[y=2x-3\]From the equation above the x - intercept is \[x=\frac{ 3 }{ 2 }\]

OpenStudy (anonymous):

Did u get it ?

OpenStudy (anonymous):

thanks u r awesome